OpenStudy (anonymous):

9/x + 4/(x - 3) = 6/(x - 3) What value of x makes this true?

OpenStudy (pfenn1):

\[\left( \frac 9x \right)+ \left( 4 \over x-3 \right)= \left( 6 \over x-3 \right)\]\[\left( \frac9x \right)=\left( 2 \over x-3 \right)\]\[9(x-3)=2x\]\[7x=3\]\[x=\frac 37\]